Session: May 1998 (session 2)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (brand at Sun May 31 07:33:02 METDST 1998) Success

Torun (K. Borkowski at Sun May 31 13:44:28 METDST 1998) Failure
No observations due to a failure of the telescope elevation drive.

Onsala (A. Polatidis at Sun May 31 14:04:25 METDST 1998) Success
No known problems

Westerbork (Tony Foley at Sun May 31 20:50:03 METDST 1998) Minor failures
Decoder still not working. Tape ran off the end of the reel (>9000ft) at 04:28UT.

Jodrell (Alastair Gunn at Mon Jun 1 15:42:58 METDST 1998) Success
No known problems.

Effelsberg (dgraham at Tue Jun 2 18:38:22 METDST 1998) Minor failures
off source 0238-0241, some rain at 0200. Program source is about 50mJy today.

Calibration (Jean\-Francois Desmurs at Mon Jun 8 13:25:22 METDST 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the may98 directory under the name ef005a.antab. Read the comments in the file to get further information.

Noto (C.Trigilio at Mon Jun 8 17:23:25 METDST 1998) Success
The field system reports some communication trouble with BBC03, but it seems to work fine.